OpenStudy (anonymous):

I need help with multiplying a radical expression. srqt14x * sqrt98x

OpenStudy (whpalmer4):

\[\sqrt{14x}\sqrt{98x}=\sqrt{14*98*x*x}=\]

OpenStudy (whpalmer4):

For any positive value of \(u\), \(\sqrt{u^2}=\sqrt{u*u}=u\)

OpenStudy (anonymous):

Thank you. I don't know what I was trying to do there. I got 14x sqrt7

OpenStudy (whpalmer4):

That is correct if x>0, good job!